Circles A, B, and C are externally tangent to each other. AB= 10 cm, BC= 12cm, and CA=16. Find the measures of the radii of circles A, B, and C.
C>A>B

Answer by Alan3354(69443)

Circles A, B, and C are externally tangent to each other. AB= 10 cm, BC= 12cm, and CA=16. Find the measures of the radii of circles A, B, and C.
C>A>B
----------
A + B = 10 --> B = -A + 10
B + C = 12
A + C = 16
-----------
Sub for B in 2nd eqn
-A + C = 2
A + C = 16
------------ Add
2C = 18
C = 9
------
A = 7
------
B = 3